OpenStudy (anonymous):

Please help. Find a formula that is used in sports or leisure activities. Be sure that that formula has at least two variables. Rewrite your formula to solve for the other variable. Show all steps in your work.

OpenStudy (ledah):

I wonder if this is even possible.... I'm a bit confused? any other directions?

OpenStudy (anonymous):

No there isn't that's why im so confused.

OpenStudy (ledah):

.... I declare a slap to the face of whoever is making you do this...

OpenStudy (anonymous):

IKR!! It's so frustrating.

OpenStudy (ledah):

I wouldnt stress this seeing as it doesnt even make sense. Don't even waste your time with a problem like this....

OpenStudy (anonymous):

I guess I'll just skip it and move on to the next. Although the next question doesn't make sense neither.

OpenStudy (ledah):

really? sheesh what in the world...

OpenStudy (anonymous):

Well thanks for trying to help :/

OpenStudy (ledah):

I even looked this up.... The internet backhanded me...

OpenStudy (anonymous):

How is it not possible? It's sort of obvious. There are many formulas used in sports. For instance, let's take F = mass * acceleration. Now, can solve for any such variable (where mass = m, a = acceleration): F / a= ma / a m = F/a F/m = ma/m F/m = a Obviously.

OpenStudy (anonymous):

I guess I'm dumb because that's no sense to me....

OpenStudy (ledah):

errr... its obvious?

OpenStudy (ledah):

for future reference.... let's not be so blunt...

OpenStudy (anonymous):

OK, I see. I'll explain in a bit more detail. My apologies for saying it's obvious. Alright, the question asks to solve for another variable in an equation. That's essentially what it's asking us to do. So, the formula F = ma is an equation used in many areas of mathematics applied to leisurely activities. This is saying force = mass * acceleration. As we would isolate a variable in a general algebraic equation, we can just manipulate all of the variables in the equation to try get other variables. Let's try solving for mass (m): \[F = ma\] We're dividing both sides of the equation by a, because we're trying to isolate m. Recall that ma/a = m. The process: \[\frac{F}{a} = \frac{ma}{a} \\ \frac{F}{a} = m \] Now, we've isolated m. That was our goal. See if you can do the same thing for a.

OpenStudy (anonymous):

Thanks @bloopman

OpenStudy (ledah):

Mind Blown.

OpenStudy (anonymous):

SAME @Ledah

OpenStudy (ledah):

If i can help on the next one i will.

OpenStudy (anonymous):

ok thanks (: @Ledah